OpenStudy (summersnow8):

IBM has a computer it calls the Blue Gene/L that can do 136.8 teracalculations per second. How many calculations can it do in a microsecond? The answer has to be in four significant digits.....

OpenStudy (anonymous):

Just do a proportion

OpenStudy (summersnow8):

well i now micro is 10^-6..... and tera is 10^12, but Im not sure what to multiply 136.8 by

OpenStudy (anonymous):

we are given (136.8) (10^12) calculations / second . we also know that 1 second = 10^6 microseconds . thus (136.8) (10^12) calculations / second * (1second / 10^6 microseconds) = (136.8) (10^6) calculations / microsecond or 136.8 mega-calculations / microsecond .

OpenStudy (summersnow8):

but i don't want mega...... i need microseconds

OpenStudy (anonymous):

My bad, 0.1368 TeraCalculations/microsecond

OpenStudy (whpalmer4):

136.8 tera calculations / 1 second = 136.8 * 10^12 calculations / second in 1 microsecond, it does 136.8 * 10^12 calculations / second * 1 second / 10^6 microseconds = 136.8 * 10^6 calculations / microsecond

OpenStudy (summersnow8):

im not following

OpenStudy (summersnow8):

draw?

OpenStudy (anonymous):

What whpalmer4 said was not in 4 digits, his answer was 136800000

OpenStudy (summersnow8):

well i dont understand how u or him got those answers

OpenStudy (whpalmer4):

136.8 *10^6 is 4 significant figures.

OpenStudy (whpalmer4):

I did not write 136800000 (which is also 4 significant figures)

OpenStudy (anonymous):

Right, my mistake, now i am completely confused.

OpenStudy (whpalmer4):

hmm, confusing everyone is not my goal :-)

OpenStudy (summersnow8):

sooo... im still confused on how you got, 136800000

OpenStudy (whpalmer4):

First, let's figure out a number for 136.8 tera calculations. 136.8 tera = 136.8 * 10^12 because tera means 10^12 Agreed so far?

OpenStudy (summersnow8):

agree

OpenStudy (whpalmer4):

Okay, it does 136.8 tera calculations per second, so that means 136.8 *10^12 calculations / second agreed?

OpenStudy (summersnow8):

yes

OpenStudy (whpalmer4):

okay, now we're just going to apply a conversion factor to change seconds to microseconds. 136.8*10^12 calc/sec * 1 sec / 10^6 microsecond = 136.8 * 10^6 calc/microsecond

OpenStudy (summersnow8):

okay, so should i keep it like that or make it 136800000

OpenStudy (whpalmer4):

in science problems you rarely go wrong by staying with scientific notation, in my experience.

OpenStudy (summersnow8):

but this is on mastering physics website..... so its on the computer, i guess i will have to see :) thanks

OpenStudy (whpalmer4):

well, as always, you have to adapt to whatever unreasonable demands are placed on you, until you don't :-)

OpenStudy (whpalmer4):

but in general, scientific notation is much more convenient when doing (wait for it) science. for example, I helped with a problem last night where the goal was to find the frequency of some violet light. I think you'll agree that 5.77 * 10^14 Hz or 577 THz is a much more convenient answer than 577000000000000 Hz :-)

OpenStudy (anonymous):

or substitution

OpenStudy (anonymous):

ooh i had mastering physics :D I rocked at that lol.

OpenStudy (anonymous):

you can also use dimensional analisys.

OpenStudy (anonymous):

 hopefully it works and cancels out